OpenStudy (anonymous):

express in simplest form: x power of 2 + 7x + 12 divided by x - 5 times 2x - 10 divided by x + 3

OpenStudy (anonymous):

is that 'x to the power of two' -> \[x^{2}\] or the xth power of two -> \[2^{x}\]?

OpenStudy (anonymous):

the first one you typed

OpenStudy (anonymous):

Okay, I can't figure out how to clearly type the equation so lets work on the numerator. Factoring, such that you have an (x+a)*(x+b)...do you know what that is?

OpenStudy (anonymous):

for this factorization, you should have (a+b)=7 and (a*b)=12.

OpenStudy (anonymous):

I am doing a worksheet and the answer the teacher gave was 2(x+4) but I don't know how to get it.

OpenStudy (anonymous):

take it one step at a time. How do you factor the numerator?

OpenStudy (anonymous):

I am very weak at algebra and have missed a lot of school because of illnesses and I don't know what I'm doing I don't know how to factor the numerator.

OpenStudy (anonymous):

I mentioned it above. do you know how to expand an equation like this: (x+a)*(x+b)?

OpenStudy (anonymous):

that is, do you know how to do the multiplication?

OpenStudy (anonymous):

Do I do it like this: xx + 7x + 12 divided by x -5 and to the other do 2x - 10 divided by x + 3?

OpenStudy (anonymous):

The FOIL method stands for First, Outside, Inside, Last; this is the order in which you multiply the numbers. Which means First: x*x = x^2 Outside: x*b Inside: x*a Last: a*b so, going backwards, from your numerator (the top part of your fraction, x^2+7x+12), you need an a and a b such that a+b=7 and a*b=12

OpenStudy (anonymous):

so your numerator factors to (x+3)*(x+4). Please be sure you understand why; that is fundamental to these kinds of problems.

OpenStudy (anonymous):

Then you can factor out/cancel out the (x+3)

OpenStudy (anonymous):

Next, note that 2x-10=2(x-5)

OpenStudy (anonymous):

is any of this making sense?

OpenStudy (anonymous):

No I'm completely lost. I was talking to a person earlier about a problem on the same worksheet.The problem was x to power of three divided by 2y to the power of two times 6y to the power of four divided by xy. She had me do xxx divided by 2 times yy and on the other side do 2 times 3 times yyyy divided by x times y then she had me cancel out the x's and the y's then multiply across to get 3x to the power of two y.

OpenStudy (anonymous):

Okay. I'm trying to think of how to do this. Do you understand the distribution property that I was describing earlier? That is, do you know how to multiplysomething like this: (x+5)*(x+1) If so, tell me what it is.

OpenStudy (anonymous):

if not, do you understand how to multiply something like this: 3*(x+2) If so tell me what it is.

OpenStudy (anonymous):

3x +6

OpenStudy (anonymous):

Okay. You don't know how to do the first one? That's okay, I just want to understand where you are. (x+5)*(x+1)